OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23 -1 1 -4 <-- this is what i got I'm not sure if its correct 4 please correct me if I'm wrong

OpenStudy (anonymous):

5x + 3y = 7 ------(1 3x - 5y = -23 ------(2 Multiplying eq (1 by by and eq (2 by 3 and adding (1 and (2 we find 25x + 15y = 35 ------(1 9x - 15y = -69 ------(2 --------------------------- 34 x= -34 x= -34/34 i.e. x= -1 Plugging x= -1 in eq (1 we find; 5*(-1) + 3y = 7 i.e. -5+3y = 7 i.e. 3y = 7+5 i.e. 3y 12 i.e. y= 12/3 i.e. y= 4 Hence the value of y is 4. @NadiaElhaj_1

OpenStudy (anonymous):

yes, you are correct :)

OpenStudy (anonymous):

thank you so much @dpasingh and @ellaponting